# Optimize Incompatible Parameters through Compatibility-aware Knowledge Integration

### v1

@@ Abstract
Deep neural networks have become foundational to advancements in multiple domains, including recommendation systems, natural language processing, and so on. Despite their successes, these models often contain incompatible parameters that can be underutilized or detrimental to model performance, particularly when faced with specific, varying data distributions. Existing research excels in removing such parameters or merging the outputs of multiple different pretrained models. However, the former focuses on efficiency rather than performance, while the latter requires several times more computing and storage resources to support inference. In this paper, we set the goal to explicitly improve these incompatible parameters by leveraging the complementary strengths of different models, thereby directly enhancing the models without any additional parameters. Specifically, we propose Compatibility-aware Knowledge Integration (CKI), which consists of Parameter Compatibility Assessment and Parameter Splicing, which are used to evaluate the knowledge content of multiple models and integrate the knowledge into one model, respectively. The integrated model can be used directly for inference or for further fine-tuning. We conduct extensive experiments on various datasets for recommendation and language tasks, and the results show that Compatibility-aware Knowledge Integration can effectively optimize incompatible parameters under multiple tasks and settings to break through the training limit of the original model without increasing the inference cost.
@@ Introduction
Figure 1: (a) shows the Incompatible Parameter issue. (b) describes Model Pruning, which removes incompatible parameters from $M_{A}$. (c) presents output ensemble, which combines the inference results of $M_{A}$ and $M_{B}$ for a final result. (d) introduces Compatibility-aware Knowledge Integration, which evaluates each parameter’s compatibility in global and local views, then integrates the knowledge of $M_{A}$ and $M_{B}$ to get model $M_{C}$. (e) shows that Compatibility-aware Knowledge Integration outperforms baselines in different scenarios.
@@
Specifically, we design a novel Unified Compatibility-aware Knowledge Integration algorithm, to address the aforementioned limitations. The core idea of our approach, is to assess the compatibility at each parameter position of each model, and then fuse the parameter across multiple models based on these uncertainties. As shown in Figure 1, Compatibility-aware Knowledge Integration comprises the following parts: (1) Parameter Compatibility Assessment includes two aspects: local-level parameter uncertainty to assess the uncertainty of each parameter position, and global-level model information content to assess the significance of the entire model. Subsequently, Parameter Compatibility Assessment fuses the local-level model uncertainty and global-level model information content into a comprehensive parameter compatibility. This dual-perspective framework is predicated on the understanding that the fidelity of model parameters is not uniformly distributed across the model’s architecture; discrepancies can arise both at the granularity of individual parameters and at the holistic level of the model’s overall parameter configuration. This distinction underpins our methodology, which synergistically integrates these two evaluative lenses into a unified parameter compatibility metric. Local-level Uncertainty Assessment examines the variance in parameter values across analogous positions within different model instances. This analysis hinges on the premise that the consistency—or lack thereof—in parameter values at a given position serves as an indicator of the reliability and uncertainty of those parameters. For example, minimal variance among models in a specific parameter suggests a convergence towards a stable, and presumably optimal, value, whereas an outlier may signal a deviation from this collective wisdom. Global-level Information Content Assessment converts model parameters into histogram distribution, and then further obtains information entropy, which is also the basis for measuring the global-level information content of the model or the global-level information content measurement. We posit that a model’s information entropy—a measure of the diversity and richness of information encapsulated within—correlates with the significance of its parameters. A model distinguished by high information entropy is posited to possess a more robust and nuanced parameter set, capable of capturing a broader spectrum of features and nuances within the training data. Dual-Perspective Compatibility Blend adopts a holistic evaluation of parameter compatibility that encapsulates both individual parameter efficacy and the collective synergy of the model’s parameters. This comprehensive approach enables a more nuanced understanding of model compatibility, facilitating the identification of models that not only perform well in aggregate but are also constructed from high-compatibility components at every level of their architecture. (2) Parameter Splicing involves splicing the parameters of multiple models based on the parameter compatibility. The information entropy of multiple models at each position is extracted nonlinear features by neural networks to be used as splicing weights. To splice the parameters, we design hard splicing and soft splicing methods. Hard Splicing adopts a decisive approach, selecting for each parameter position the most optimal parameter from the pool of pre-trained models based on compatibility assessment. This method ensures that the spliced model is made up of the best parts available, albeit with the caveat of potential information loss due to the outright exclusion of lower-compatibility parameters. To mitigate this, we introduce Soft Splicing, which calculates the weights of the spliced model under the guidance of the weighted sum of parameter compatibility and parameter weights, maximizing the preservation of parameter information and optimizing model parameters.
@@
It should be noted that, the integrated model has the same structure and the same number of parameters as the original pre-trained models, so it does not add even a little bit of additional inference cost. Our Compatibility-aware Knowledge Integration algorithm has a very wide range of applications. It can optimize parameters that are incompatible with the global data distribution as well as those that are incompatible with the individual data distribution. The integrated model can be used for inference without any retraining. Furthermore, the integrated model can serve as a better initialization model, requiring only one epoch of retraining to achieve better performance.
@@
  • We instantiated the incompatible parameter optimization method and designed a universal Compatibility-aware Knowledge Integration framework consists of Parameter Compatibility Assessment and Parameter Splicing.
@@
Incompatible Parameters. Deep learning models often suffer from excessive parameters and redundancies. Network pruning (Guo, Ouyang, and Xu 2020; Wang et al. 2020; Sanh, Wolf, and Rush 2020; Han et al. 2015; Dong, Chen, and Pan 2017; Ma, Fang, and Wang 2023; Zhou et al. 2023; Fang, Ma, and Wang 2024) aims to eliminate non-essential parameters to maintain effectiveness. Evaluation methods include second-order derivatives of the loss function, parameter magnitudes, and other metrics. Output Ensemble (Dong et al. 2020; Lee et al. 2021; Zhou et al. 2021; Kumar et al. 2023; Mazari, Boudoukhani, and Djeffal 2024; Yuan, Li, and Hao 2023) aggregates results from multiple models trained under different conditions, optimizing final output without altering model parameters.
@@
Deep Learning Models. Deep learning models have achieved excellent performance in various tasks. In vision tasks, models such as ResNet (He et al. 2016), SqueezeNet (Iandola et al. 2016), ShuffleNet (Zhang et al. 2018; Ma et al. 2018), MobileNet (Howard et al. 2017; Sandler et al. 2018; Howard et al. 2019) have significantly improved efficiency. In language tasks, GPT (Radford et al. 2018, 2019; Brown et al. 2020), BERT (Devlin et al. 2018), and Roberta (Liu et al. 2019; Devlin et al. 2018) excel through pre-training and attention mechanisms. In recommendation tasks, models like DeepFM (Guo et al. 2017), LightGCN (He et al. 2020), DIN (Zhou et al. 2018), GRU4Rec (Hidasi et al. 2016), SASRec (Kang and McAuley 2018), Bert4Rec (Sun et al. 2019), and ICLRec (Chen et al. 2022) are often used. Dynamic neural networks are now widely used in computer vision (Alaluf et al. 2022; Dinh et al. 2022), recommendation systems (Lv et al. 2024a; Yan et al. 2022), large language models (Zhang et al. 2024), device-cloud collaboration (Lv et al. 2023, 2024b), and other fields. They can adjust model parameters based on samples, enabling model personalization.
@@
Figure 2: Overview of the proposed Compatibility-aware Knowledge Integration. Our Compatibility-aware Knowledge Integration includes two parts: Parameter Compatibility Assessment and Parameter Splicing. (a) describes the Parameter Compatibility Assessment. It consists of 3 parts: (a1) Local-level Parameter Uncertainty Assessment, (a2) Global-level Model Information Content Assessment, and (a3) Dual-Perspective Parameter Compatibility Assessment. (b) describes the Parameter Splicing, which includes (b1) Hard Splicing and (b2) Soft Splicing. (c) describes the extension of Compatibility-aware Knowledge Integration from 2 models to multiple models.
Compatibility-aware Knowledge Integration. CKI also assumes that based on the dataset $\mathcal{D}_{\rm{Train}}$ like Output Ensemble, $n$ models are trained (where $n\geq 2$). For simplicity, let’s initially set $n$ to 2 and assume that the parameters of the two trained models are $W_{A}$ and $W_{B}$. Similar to Model Pruning, let’s assume that the uncertainty matrices for $W_{A}$ and $W_{B}$ are $V^{(A)}$ and $V^{(B)}$, respectively. Then, CKI can be formalized as,
@@
Compatibility-aware Knowledge Integration for Dual Models
In this section, we will introduce our Compatibility-aware Knowledge Integration. As shown in Figure 2, our Compatibility-aware Knowledge Integration includes two parts: Parameter Compatibility Assessment and Parameter Splicing. (a) Describes the Parameter Compatibility Assessment. It consists of three parts: (a1) describes the Local-level Parameter Uncertainty Assessment method, which is used to measure the uncertainty of each parameter in each model. (a2) describes the Global-level Model Information Content Assessment, which is used to measure the model information content from a global view. (a3) describes the Dual-Perspective Parameter Compatibility Assessment, which assesses model uncertainty from both local-level and global-level perspectives. (b) Describes the Parameter Splicing, which includes two parts: (b1) describes the hard splicing, where low-compatibility parameters at each position directly replace high-compatibility parameters from two models. (b2) describes the soft splicing, where each parameter is weighted and fused based on the compatibility of the parameter at that position in the two models. (c) describes the extension of Compatibility-aware Knowledge Integration from 2 models to multiple models. The details are elaborated in the following subsections.
@@
Generalization for Compatibility-aware Knowledge Integration. In order to better understand the generalization of Compatibility-aware Knowledge Integration, we take hard splicing as an example and use classical generalization theory to analyze the generalization boundary of the spliced model. We show that under some assumptions, the generalization boundary of spliced model $W$ can be bounded, see Appendix for more theoretical information.
Compatibility-aware Knowledge Integration for Multiple Models
@@
Table 1: Performance comparison of the proposed method and baselines on recommendation tasks when the pre-trained model is a static models. The best performance is highlighted in bold while the second best performance is underlined.
Dataset Methods Metric
@@
Table 2: Performance comparison of the proposed method and baselines on language tasks. The best performance is highlighted in bold while the second best performance is underlined.
@@
Table 3: Performance comparison of the proposed method and baselines on recommendation task when pre-trained models include both static and dynamic models. The best value is highlighted in bold while the second best value is underlined.
@@
Compatibility-aware Knowledge Integration for Global Incompatible Parameters
@@
Compatibility-aware Knowledge Integration for Individual Incompatible Parameters
Table 3 compares the performance of parameter concatenation based on the dynamic recommendation models (Model $M_{B}$ (G)). Unlike the parameter concatenation method between static models, the parameters of the Model $M_{B}$ (G) are generated based on the sample using a HyperNetwork. The dynamic model and static model use the same base model for parameter concatenation. From the table, we have the following findings: (1) across all datasets and metrics, the performance of parameter concatenation significantly surpasses other baselines, for instance, improving from the best baseline’s NDCG@5=0.2988 to NDCG@5=0.4272, and HR@5=0.4272 to HR@5=0.4592. (2) Comparing Figure 3 and 4, we find that the improvement brought about by the parameter concatenation of static and dynamic models over the baseline is more significant than the improvement brought by parameter concatenation among static models. This is because the model parameters of the dynamic model are customized based on the input samples in an unsupervised process. Therefore, although the performance of the dynamic model is far superior to that of the static model, it also has greater instability. These parameters can be compensated for with a static model. Experimental results show that our method significantly improves model performance. This indicates that our approach is effective in optimizing parameters that are incompatible with the individual data distribution.
Compatibility-aware Knowledge Integration for Initialization and Re-training
@@
Table 4: Compatibility-aware Knowledge Integration for model initialization and re-training. The best performance is highlighted in bold.
After the model parameters are integrated, the resulting model can be directly used for inference, or it can serve as a better initial parameter set, allowing the model to be further trained on this basis for improved results. To verify the initial performance of the model after Compatibility-aware Knowledge Integration, we trained the model again post-integrating. As shown in Table 4, We found that the integrated model only needs one epoch to converge and achieve better results than the parameter integrated model and other baselines. This also demonstrates the benefits of Compatibility-aware Knowledge Integration for model initialization.
@@ Conclusion
We introduced Compatibility-aware Knowledge Integration, a novel method to optimize incompatible parameters in deep neural networks by leveraging the complementary strengths of different pretrained models, without adding extra parameters. Our experiments on recommendation and language tasks demonstrate that Compatibility-aware Knowledge Integration effectively enhances model performance without increasing inference costs, offering a promising approach for deploying more robust and efficient models.
ACKNOWLEDGEMENTS

### v2

@@ Abstract
Deep neural networks have become foundational to advancements in multiple domains, including recommendation systems, natural language processing, and so on. Despite their successes, these models often contain incompatible parameters that can be underutilized or detrimental to model performance, particularly when faced with specific, varying data distributions. Existing research excels in removing such parameters or merging the outputs of multiple different pretrained models. However, the former focuses on efficiency rather than performance, while the latter requires several times more computing and storage resources to support inference. In this paper, we set the goal to explicitly improve these incompatible parameters by leveraging the complementary strengths of different models, thereby directly enhancing the models without any additional parameters. Specifically, we propose Compatibility-aware Knowledge Integration (CKI), which consists of Parameter Compatibility Assessment and Parameter Splicing, which are used to evaluate the knowledge content of multiple models and integrate the knowledge into one model, respectively. The integrated model can be used directly for inference or for further fine-tuning. Extensive experiments on various recommendation and language datasets show that CKI can effectively optimize incompatible parameters under multiple tasks and settings to break through the training limit of the original model without increasing the inference cost.
@@ Introduction
Figure 1: (a) shows the Incompatible Parameter issue. (b) describes Model Pruning, which removes incompatible parameters from $M_{A}$. (c) presents output ensemble, which combines the inference results of $M_{A}$ and $M_{B}$ for a final result. (d) introduces CKI, which evaluates each parameter’s compatibility in global and local views, then integrates the knowledge of $M_{A}$ and $M_{B}$ to get model $M_{C}$. (e) shows that CKI outperforms baselines in different scenarios.
@@
Specifically, we design a novel Unified Compatibility-aware Knowledge Integration (CKI) algorithm, to address the aforementioned limitations. The core idea of our approach, is to assess the compatibility at each parameter position of each model, and then fuse the parameter across multiple models based on these uncertainties. As shown in Figure 1, CKI comprises the following parts: (1) Parameter Compatibility Assessment includes two aspects: local-level parameter uncertainty to assess the uncertainty of each parameter position, and global-level model information content to assess the significance of the entire model. Subsequently, Parameter Compatibility Assessment fuses the local-level model uncertainty and global-level model information content into a comprehensive parameter compatibility. This dual-perspective framework is predicated on the understanding that the fidelity of model parameters is not uniformly distributed across the model’s architecture; discrepancies can arise both at the granularity of individual parameters and at the holistic level of the model’s overall parameter configuration. This distinction underpins our methodology, which synergistically integrates these two evaluative lenses into a unified parameter compatibility metric. Local-level Uncertainty Assessment examines the variance in parameter values across analogous positions within different model instances. This analysis hinges on the premise that the consistency—or lack thereof—in parameter values at a given position serves as an indicator of the reliability and uncertainty of those parameters. For example, minimal variance among models in a specific parameter suggests a convergence towards a stable, and presumably optimal, value, whereas an outlier may signal a deviation from this collective wisdom. Global-level Information Content Assessment converts model parameters into histogram distribution, and then further obtains information entropy, which is also the basis for measuring the global-level information content of the model or the global-level information content measurement. We posit that a model’s information entropy—a measure of the diversity and richness of information encapsulated within—correlates with the significance of its parameters. A model distinguished by high information entropy is posited to possess a more robust and nuanced parameter set, capable of capturing a broader spectrum of features and nuances within the training data. Dual-Perspective Compatibility Blend adopts a holistic evaluation of parameter compatibility that encapsulates both individual parameter efficacy and the collective synergy of the model’s parameters. This comprehensive approach enables a more nuanced understanding of model compatibility, facilitating the identification of models that not only perform well in aggregate but are also constructed from high-compatibility components at every level of their architecture. (2) Parameter Splicing involves splicing the parameters of multiple models based on the parameter compatibility. The information entropy of multiple models at each position is extracted nonlinear features by neural networks to be used as splicing weights. To splice the parameters, we design hard splicing and soft splicing methods. Hard Splicing adopts a decisive approach, selecting for each parameter position the most optimal parameter from the pool of pre-trained models based on compatibility assessment. This method ensures that the spliced model is made up of the best parts available, albeit with the caveat of potential information loss due to the outright exclusion of lower-compatibility parameters. To mitigate this, we introduce Soft Splicing, which calculates the weights of the spliced model under the guidance of the weighted sum of parameter compatibility and parameter weights, maximizing the preservation of parameter information and optimizing model parameters.
@@
It should be noted that, the integrated model has the same structure and the same number of parameters as the original pre-trained models, so it does not add even a little bit of additional inference cost. Our CKI algorithm has a very wide range of applications. It can optimize parameters that are incompatible with the global data distribution as well as those that are incompatible with the individual data distribution. The integrated model can be used for inference without any retraining. Furthermore, the integrated model can serve as a better initialization model, requiring only one epoch of retraining to achieve better performance.
@@
  • We instantiated the incompatible parameter optimization method and designed a universal CKI framework consists of Parameter Compatibility Assessment and Parameter Splicing.
@@
Incompatible Parameters. Deep learning models often suffer from excessive parameters and redundancies. Network pruning (Guo, Ouyang, and Xu 2020; Wang et al. 2020; Sanh, Wolf, and Rush 2020; Han et al. 2015; Dong, Chen, and Pan 2017; Ma, Fang, and Wang 2023; Zhou et al. 2023; Fang, Ma, and Wang 2024) aims to eliminate non-essential parameters to maintain effectiveness. Evaluation methods include second-order derivatives of the loss function, parameter magnitudes, and other metrics. Output Ensemble (Lee et al. 2021; Zhou et al. 2021; Kumar et al. 2023; Mazari, Boudoukhani, and Djeffal 2024; Yuan, Li, and Hao 2023; Zhu et al. 2024; Wang et al. 2024) aggregates results from multiple models trained under different conditions, optimizing final output without altering model parameters.
@@
Deep Learning Models. Deep learning models have achieved excellent performance in various tasks. In vision tasks, models such as ResNet (He et al. 2016), SqueezeNet (Iandola et al. 2016), MobileNet (Howard et al. 2017; Sandler et al. 2018; Howard et al. 2019) have significantly improved efficiency. In language tasks, GPT (Radford et al. 2018, 2019; Brown et al. 2020), BERT (Devlin et al. 2018), and Roberta (Liu et al. 2019; Devlin et al. 2018) excel through pre-training and attention mechanisms. In recommendation tasks, models like DeepFM (Guo et al. 2017), LightGCN (He et al. 2020), DIN (Zhou et al. 2018), GRU4Rec (Hidasi et al. 2016), SASRec (Kang and McAuley 2018), Bert4Rec (Sun et al. 2019), and ICLRec (Chen et al. 2022) are often used. Dynamic neural networks are now widely used in computer vision (Alaluf et al. 2022; Dinh et al. 2022), recommendation systems (Lv et al. 2024a; Yan et al. 2022), large language models (Zhang et al. 2024), device-cloud collaboration (Lv et al. 2023, 2024b), and other fields. They can adjust model parameters based on samples, enabling model personalization.
@@
Figure 2: Overview of the proposed CKI. Our CKI includes two parts: Parameter Compatibility Assessment and Parameter Splicing. (a) describes the Parameter Compatibility Assessment. It consists of 3 parts: (a1) Local-level Parameter Uncertainty Assessment, (a2) Global-level Model Information Content Assessment, and (a3) Dual-Perspective Parameter Compatibility Assessment. (b) describes the Parameter Splicing, which includes (b1) Hard Splicing and (b2) Soft Splicing. (c) describes the extension of CKI from 2 models to multiple models.
CKI. CKI also assumes that based on the dataset $\mathcal{D}_{\rm{Train}}$ like Output Ensemble, $n$ models are trained (where $n\geq 2$). For simplicity, let’s initially set $n$ to 2 and assume that the parameters of the two trained models are $W_{A}$ and $W_{B}$. Similar to Model Pruning, let’s assume that the uncertainty matrices for $W_{A}$ and $W_{B}$ are $V^{(A)}$ and $V^{(B)}$, respectively. Then, CKI can be formalized as,
@@
CKI for Dual Models
In this section, we will introduce our CKI. As shown in Figure 2, our CKI includes two parts: Parameter Compatibility Assessment and Parameter Splicing. (a) Describes the Parameter Compatibility Assessment. It consists of three parts: (a1) describes the Local-level Parameter Uncertainty Assessment method, which is used to measure the uncertainty of each parameter in each model. (a2) describes the Global-level Model Information Content Assessment, which is used to measure the model information content from a global view. (a3) describes the Dual-Perspective Parameter Compatibility Assessment, which assesses model uncertainty from both local-level and global-level perspectives. (b) Describes the Parameter Splicing, which includes two parts: (b1) describes the hard splicing, where low-compatibility parameters at each position directly replace high-compatibility parameters from two models. (b2) describes the soft splicing, where each parameter is weighted and fused based on the compatibility of the parameter at that position in the two models. (c) describes the extension of CKI from 2 models to multiple models. The details are elaborated in the following subsections.
@@
Generalization for CKI. In order to better understand the generalization of CKI, we take hard splicing as an example and use classical generalization theory to analyze the generalization boundary of the spliced model. We show that under some assumptions, the generalization boundary of spliced model $W$ can be bounded, see Appendix for more theoretical information.
CKI for Multiple Models
@@
Table 1: Performance comparison on recommendation tasks when the pre-trained model is a static models.
Datasets Methods Metric
@@
Table 2: Performance comparison on language tasks.
@@
Table 3: Performance comparison on recommendation task when pre-trained models include both static and dynamic models.
@@
To facilitate performance comparison, the best value is highlighted in bold, while the second best value is underlined. Some tables only highlight the maximum value.
CKI for Global Incompatible Parameters
@@
CKI for Individual Incompatible Parameters
Table 3 compares the performance of parameter fusion based on dynamic recommendation models (model $M_{B}$ (G)). Unlike the parameter fusion methods for static models, the parameters of $M_{B}$ (G) are dynamically generated by a hypernetwork based on the input samples. Both dynamic and static models use the same base model for parameter fusion. Across all datasets and metrics, parameter fusion significantly outperforms other baselines. Additionally, Comparing Table 1 and Table 3, we find that parameter fusion between dynamic and static models achieves a more significant performance improvement over baseline models than parameter fusion between static models. This is because the parameters of dynamic models are adaptively generated in an unsupervised manner based on input samples. Although the performance of dynamic models significantly surpasses that of static models, they also exhibit greater instability. This instability can be compensated by incorporating parameters from static models. Experimental results demonstrate that this method is highly effective in optimizing parameters that are incompatible with individual data distributions.
CKI for Initialization and Re-training
@@
Table 4: Compatibility-aware Knowledge Integration for model initialization and re-training.
After the model parameters are integrated, the resulting model can be directly used for inference, or it can serve as a better initial parameter set, allowing the model to be further trained on this basis for improved results. To verify the initial performance of the model after CKI, we trained the model again post-integrating. As shown in Table 4, We found that the integrated model only needs one epoch to converge and achieve better results than the parameter integrated model and other baselines. This also demonstrates the benefits of CKI for model initialization.
@@ Conclusion
We introduced CKI, a novel method to optimize incompatible parameters in deep neural networks by leveraging the complementary strengths of different pretrained models, without adding extra parameters. Our experiments on recommendation and language tasks demonstrate that CKI effectively enhances model performance without increasing inference costs, offering a promising approach for deploying more robust and efficient models.
Acknowledgments
@@
Appendix A Appendix
Supplementary Methodology
Pseudo Code
Algorithm 1 shows the pseudo code of Compatibility-aware Knowledge Integration.
Algorithm 1 Optimize Incompatible Parameters through Compatibility-aware Knowledge Integration
Variable: Data $\mathcal{X}=\{x\}$, Local-level Parameter Uncertainty Set $\mathcal{V_{L}}=\{V_{L}\mid M_{k}\in\mathcal{M}\}$, Global-level Model Information Content Set $\mathcal{V_{G}}=\{V_{G}\mid M_{k}\in\mathcal{M}\}$, Parameter Compatibility Set $\mathcal{V}=\{V\mid M_{k}\in\mathcal{M}\}$, Pretrained Model Set $\mathcal{M}$, Integrated Model $M$, Local-level Parameter Uncertainty Network $f_{L}$, Global-level Model Information Content Network $f_{G}$.
[…]
Repeat
 If $f_{L}$ and $f_{G}$ have not yet been well-trained
 I. Do forward propagation to get $\mathcal{V_{L}}$ as Eq.14, $\mathcal{V_{G}}$ as Eq.15.
 II. Choose a splicing method from hard splicing Eq.11 and soft splicing Eq.12.
 III. Train as $\mathop{\rm min}_{f_{L},f_{G}}\mathcal{L}=\sum\nolimits_{x\in\mathcal{X},y\in\mathcal{Y}}l_{\text{}}(y,\hat{y}=M(x))$
Until Convergence
Training and Inference Procedure
Training Procedure. During training, in the dual-model CKI, what we need to do is obtain the optimal $V=\{V^{(A)},V^{(B)}\}$ to minimize the loss $\mathcal{L}$. When extending to multi-model CKI, our goal is similar. Assuming the number of models is $N$, the optimization objective is to find the optimal compatibility set $V=\{V^{(k)}\}_{k=1}^{n}$ to minimize $\mathcal{L}$, that is, $\operatorname*{arg\,min}_{V=\{V^{(k)}\}_{k=1}^{n}}\mathcal{L}$.
Inference Procedure. After the splicing of model parameters, the parameters $W=\{W_{k}\}_{k=1}^{n}$ of $n$ models are combined into the parameters of one model $W$, and then this model can perform inference just like a regular recommendation model, that is, $\hat{y}=Wx$.
Assessment Methods Metrics
Local Uncertainty Global Information Content AUC UAUC NDCG@5 HR@5
0.8557 0.8534 0.3779 0.5036
0.8554 0.8534 0.3797 0.5057
0.8582 0.8558 0.3803 0.5064
0.8585 0.8560 0.3823 0.5071
Table 5: Results of the ablation studies w.r.t. the assessment methods. The best performance is highlighted in bold while the second best performance is underlined.
Supplementary Experiments
Evaluation Metrics
We use Cost to measure the model’s consumption of resources. The calculation of Cost is the inference time of the method divided by the inference time of the base model, rounded to one decimal place.
The Effect of Parameter Compatibility Assessment Methods
[…]
The Effect of Parameter Splicing Methods
Parameter Splicing is including hard splicing and soft splicing. Table 6, shows the comparison of hard splicing and soft splicing. The experimental results indicate that soft splicing is a superior method compared to hard splicing.
Fusion Methods Metrics
AUC UAUC NDCG@5 HR@5
Hard splicing 0.8561 0.8532 0.3806 0.5047
Soft splicing 0.8581 0.8558 0.3805 0.5064
Table 6: Results of the ablation studies w.r.t. the splicing methods. The best performance is highlighted.
Metrics Metrics
Datasets Base Models Methods NDCG@5 HR@5 NDCG@10 HR@10 Datasets Base Models Methods NDCG@5 HR@5 NDCG@10 HR@10
Model A 0.4027 0.5260 0.4424 0.6490 Model A 0.3779 0.5036 0.4179 0.6273
Model B 0.4011 0.5255 0.4428 0.6544 Model B 0.3763 0.5034 0.4173 0.6301
DIN Ours 0.4095 0.5344 0.4505 0.6612 DIN Ours 0.3805 0.5064 0.4212 0.6326
Model A 0.4094 0.5329 0.4525 0.6658 Model A 0.3677 0.4988 0.4109 0.6321
Model B 0.4059 0.5303 0.4481 0.6602 Model B 0.3718 0.5024 0.4150 0.6360
GRU4Rec Ours 0.4100 0.5350 0.4527 0.6669 GRU4Rec Ours 0.3756 0.5091 0.4162 0.6347
Model A 0.4033 0.5262 0.4445 0.6548 Model A 0.3734 0.5038 0.4160 0.6358
Model B 0.4031 0.5266 0.4451 0.6565 Model B 0.3772 0.5090 0.4182 0.6357
Books SASRec Ours 0.4043 0.5276 0.4468 0.6573 Music SASRec Ours 0.3799 0.5111 0.4221 0.6416
Table 7: The impact of choosing different small models on performance.
Figure 3 and 4 shows the comparison of our method with baselines on UAUC, AUC, and Cost. Each bar in the figure represents UAUC or AUC, while the line represents Cost. All these tables and figures demonstrate that our method improved performance without increasing the cost.
The Impact of the Different Base Models
[…]
Figure 5: The impact of the number of integrated models.
The Impact of the Number of Integrated Models
As shown in Figure 5, we expanded model integrating from dual models to multiple models. The experimental results show: the best performance is achieved when the number of models is 7. When the number of models is less than 7, the performance of the integrated model increases with the increase in the number of pretrained models. This is because, within a certain range, integrating of multiple model can more effectively leverage the knowledge learned from other models, optimizing the redundant parameters of the current model. However, an increase in the number of pretrained models leads to an increase in the size of the integrated network, and the available data volume can no longer support the training of a larger integrated network. Therefore, when the number of models exceeds 7, the performance of the integrated model rapidly declines with the increase in the number of pretrained models.
Hyperparameters and Training Schedules
We summarize the hyperparameters and training schedules of the datasets used in the experiments in Table 8.
Dataset Parameters Setting
SST-2 RTE GPU Tesla A100
Optimizer Adam
Learning rate 0.001
Batch size 32
the Dimension of embedding 1×514
Amazon Beauty Douban Book Douban Music Movielens-1M GPU Tesla A100
Optimizer Adam
Learning rate 0.001
Batch size 1024
Sequence length 10
the Dimension of embedding 1×32
Table 8: Hyperparameters of CKI.